OpenStudy (anonymous):

What is the kinetic energy of a 45kg object moving at 13 m/sec?

OpenStudy (anonymous):

@CallMeKiki plz help!

OpenStudy (anonymous):

@CallMeKiki plzzzzz help meeee!

OpenStudy (callmekiki):

What are the options?

OpenStudy (callmekiki):

Please wait.

OpenStudy (callmekiki):

First we identify the information we are given in the problem: mass = 45 kg velocity = 13 m/sec Next, we place this information into the kinetic energy formula: KE = 1/2 mv2 KE = 1/2 (45 kg)(13 m/sec)2 Solving the equation gives a kinetic energy value of 3802.5 J

OpenStudy (anonymous):

3803 joules. Medal plz :)

OpenStudy (callmekiki):

Lol @Skai